Bud, Inc. sold 10,000 units for $70/unit of its only product last year. Operating information from...

  1. Bud, Inc. sold 10,000 units for $70/unit of its only product last year. Operating information from last year is shown below:                                                                                            

                                                                                    Total Cost            

            Total Manufacturing Costs                                 $440,000                                  

            Selling & Administrative Expenses                      

                        Variable:                                               $ 60,000

                        Fixed:                                                   $ 32,000

Bud management has indicated that manufacturing costs are 50% variable and 50% fixed. Management does not expect a change in the price/cost structure for next year.

  1. What percentage of each sale goes toward profit after the breakeven point is reached?
  2. If sales increase by $70,000, how much will net income increase?
  3. How many units must be sold to breakeven?
  4. If the company sells 8,000 units, what is expected profit?
  5. The marketing manager believes that an increase in advertising of $30,000 would increase expected sales by 1,000 units. How would this affect (1) the breakeven point and (2) profit?
Answer #1

Given Number of units of sales = 10,000

Sale price per unit = $70

Total Sales = Number of units * Sale price per unit = 10,000 * 70 = $700,000

Variable Selling & Administrative Expenses = $60,000

Variable Selling & Administrative Expenses per unit = Total Expenses / Number of units = $60,000/ 10,000 = $6 per unit

Variable manufacturing costs = $440,000 * 50% = $220,000

Variable manufacturing costs per unit = Total Expenses / Number of units = $220,000 / 10,000 = $22 per unit

Fixed Manufacturing costs = $440,000 * 50% = $220,000

Particulars Number of Units (i) Per Unit (ii) Amount (iii)=(ii)*(i)
(A) Selling Price                         10,000 $       70.00 $                  700,000
(B) Variable Expenses
(a) Manufacturing 10,000 $       22.00 $                  220,000
(b) Selling 10,000 $         6.00 $                    60,000
Contribution (C)=(A)-(B) $                  420,000
(D) Fixed Expenses
(a) Manufacturing $                  220,000
(b) Selling $                    32,000
(E) Profit [(C)-(D)] $                  168,000

Contribution per unit = $420,000/10,000 = $42

(a) Percentage of each sale going toward profit after breakeven point:

At the breakeven point, the contribution fully recovers the fixed cost. Subsequently for each sale, the contribution results in profit. Hence after the breakeven point percentage of the sale, going toward profit is contribution percentage which is calculated as below:

Contribution percentage = Contribution per unit/ Selling Price per unit = $42/$70 = 0.6 = 60%

(b) Increase in net income if sales increase by $70,000:

Since fixed costs do not change with the increase in sales, Increase in net profit is increase in contribution.

Increase in Contribution = Increase in sales * contribution percentage = $70,000 * 60% = $42,000

(c) Number of units to be sold to breakeven:

Breakeven point(in units) = Fixed Costs / Contribution per unit

Breakeven point(in units) = ($220,000+$32,000)/$42 = 6,000 units

(d) Expected profit if company sells 8,000 units:

Profit = Sales * Contribution Percentage - Fixed Costs = (8,000*$70)*60% - $252,000 = $84,000

(e) Increase in advertising by $30,000 to increase sales of 1,000 units:

Here increased costs are assumed to be fixed in nature.

New Breakeven point = Fixed Costs/ Contribution per unit=($252,000+$30,000)/$42=6,714.29 units. Hence breakeven point is increased by 714.29 units (30,000/$42)

New profit = Sales * Contribution Percentage - Fixed Costs =(11000*$70)*60%-$282,000 = $180,000. Hence profit is increased by $12,000 [(1,000*42)-30000]